Posted this before, and recieved the wrong answer. They put -3.91e6 and 2.38e7

Two point charges are on the y-axis. A 5.0 µC charge is located at y = 1.15 cm, and a -2.26 µC charge is located at y = -1.90 cm.

(a) Find the total electric potential at the origin.

(b) Find the total electric potential at the point having coordinates (1.50 cm, 0).
Answer #1

The electric potential from each charge is K*q/r, where r is the distance to the charge measured from the origin. Potentials add algebraically. For the first charge,

V1 = K*5.0*10^-6/0.0115

V1 = 0.434*10^-3*k

for the second

V2 = K*(-2.26*10^-6)/(-0.019).

V2 = 0.118*10^-3*k

The total potential at the origin is V1 + V2.

V1 + V2 = 0.552*10^-3*k

V1 + V2 = 0.552*10^-3*9*10^9

V1 + V2 = 4.968*10^6 J/c